Status Report

NOAA GOES SATOPS Morning Report 1 Feb 2001

By SpaceRef Editor
February 1, 2001
Filed under ,

SATOPS MORNING REPORT – February 1, 2001

POES

No change in the status of the POES spacecraft.

NOAA-14 rev 31384 / F at 1347Z on January 31:  L5B noisy from spacecraft;
98.1% recovered.  Further recovery attempts not possible due to spacecraft
recorder management constraints.

NOAA-15 rev 14136 / W at 1407Z on January 31:  G3B noisy from spacecraft;
99.2% recovered.  Further recovery attempts not possible due to spacecraft
recorder management constraints.

NOAA-12 rev 50465 / F at 1721Z on January 31:  G3A short from Antenna
“A” system; 96.5% data recovered.  Data recovered off Antenna “C”
system, which was shadowing the pass and retrieved all data.

NOAA-16 rev 1867 / W at 1917Z on January 31:  CDA had trouble establishing
S-band uplink due to combiner problems.  Combiner was recycled. 
Resulting losses:  L5B short (23.9%), not recoverable.  L1A short
(88.9%), no re-ship required.

To support troubleshooting of the NOAA-16 CPU/CIU timing problem, a
test sequence was commanded successfully to NOAA-16 at 12:38Z (rev 1877
/ F, 76 deg elevation) on February 1.

POES operations were nominal over the past 24 hours.

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

GOES-11 Z-Axis Precession (storage mode) re-spin #1 was successfully
executed at 0555z on February 1, 2001.  The initial spin cone angle
was nailed at 0.3 degree.  Maximum battery DOD was 3.2%.  Post
re-spin ranging starts at 0800Z.

 

GOES operations were nominal over the past 24 hours.

 

Questions or comments to msimpson@nesdis.noaa.gov

SpaceRef staff editor.